OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} x[\sin ^{-1}(\frac{ x+1 }{ 2x+1 }) - \frac{ \Pi }{ 6 } ]\]

OpenStudy (amistre64):

... Lhop rule? is my first thought

OpenStudy (anonymous):

First we have to have to indeterminate form..ryt

OpenStudy (amistre64):

yes; 0/0 format is good

OpenStudy (anonymous):

But I did Everything but i never got the answer...

OpenStudy (amistre64):

the 1/x tho might prove futile tho. since the derivatices all have and x under

OpenStudy (amistre64):

never tried it out on inf/inf tho

OpenStudy (amistre64):

\[[\sin ^{-1}(\frac{ x+1 }{ 2x+1 }) - \frac{ \Pi }{ 6 } ]'=\frac{1}{\sqrt{1-\left(\frac{ x+1 }{ 2x+1 }\right)^2}}\frac{-1}{4x^2+4x+1}\] \[[x^{-1}]'=-x^{-2}\] does that stuff simplify any?

OpenStudy (anonymous):

\[{d\over dx}x\left[\sin^{-1}\left(x+1\over 2x+1\right)-{\pi\over6}\right]\\ =\left[\sin^{-1}\left(x+1\over 2x+1\right)-{\pi\over6}\right]+x\frac{1}{\Large\sqrt{1-\left({x+1\over2x+1}\right)^2}} \]

OpenStudy (anonymous):

Aha..thxx guys i found my mistake.......